The volume of a cone shaped hole is 16 on ft cubes if the hole is 3 feet deep is hat is the radius of the hole?


1
Expert's answer
2021-11-22T10:49:09-0500

Volume =B.A*height =1/3"\Pi"rh

Thus Volume =16ft3 and Height= 3ft

Replacing in the equation

16=1/3"\Pi"r3

Reducing the numbers by canceling 1/3 with 3

so it will be:

16 = Πr2

r2=16/Π =16/3.14

taking square roots of both sides

Therefore, radius = 2.25676